OpenStudy (anonymous):

Part 1: Solve the inequality -5x + 8 less than or greater to x - 4 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (anonymous):

can someone help

OpenStudy (anonymous):

Is this your question? \(-5x+8 \leq x-4\)

OpenStudy (anonymous):

You need to isolate the x just like an equation.

OpenStudy (anonymous):

so it would be -6x + 8 \[\le\] x - 4

OpenStudy (anonymous):

No. When you subtract the x from both sides, it cancels out on the right side.

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

\[-5x-x+8 \leq \cancel{x-x}-4\] \[-6x+8 \leq -4\]

OpenStudy (anonymous):

oh so now its -6x + 8 <= -4 - 8 -8 -6x<= -12

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so it would be x <= -2

OpenStudy (anonymous):

The only difference between solving an equation and an inequality is that the inequality reverses when you multiply or divide by a negative.

OpenStudy (anonymous):

Also, remember that a negative divided by a negative has a positive answer.

OpenStudy (anonymous):

so \[x \ge -2\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

whats part 2?

OpenStudy (anonymous):

Draw the graph and then describe what it will looks like.

OpenStudy (anonymous):

This should be drawn on a number line.